OpenStudy (anonymous):

7x+3x-(10x+2)=5+x So simple but so stuck :)

OpenStudy (anonymous):

remove brackets.

OpenStudy (anonymous):

10x-10x-2=5+x -2=5+x -7=x

OpenStudy (anonymous):

ok thats what we got just looked wrong